OpenStudy (anonymous):

1. y = –x + 2 y = 3x – 1 (1 point) I need HElp

jigglypuff314 (jigglypuff314):

to solve for this system of equations you can use the substitution method y = –x + 2 y = 3x – 1 because y = y -x + 2 = 3x - 1 so solve for x this way :)

jigglypuff314 (jigglypuff314):

–4x + 3y = –12 –2x + 3y = –18 elimination method so 2x = -6 solve for x then once you get x, plug it back into on of the original equations to get y

jigglypuff314 (jigglypuff314):

x = –4y + 4 2x + 8y = 8 substitution method 2(-4y + 4) + 8y = 8 -8y + 8 + 8y = 8 the 8y's cancel out so you get 8 = 8 therefore infinite

jigglypuff314 (jigglypuff314):

y = 6x + 2 3y – 18x = 12 substitution method 3(6x + 2) - 18x = 12 the 18x's cancel out 6 does not equal 12 so no solution

jigglypuff314 (jigglypuff314):

x – 2y = 6 3x – 6y = 18 if you try elimination method you'll get the same equations so infinite

jigglypuff314 (jigglypuff314):

y – 5x = –6 3y – 15x = –12 if you try elimination method you'll get same variable terms but a different last term so no solution

jigglypuff314 (jigglypuff314):

x ≥ 4 closed point at 4 shaded to the right.

jigglypuff314 (jigglypuff314):

y < –2 open point at y = -2 shaded below

jigglypuff314 (jigglypuff314):

A local citizen wants to fence a rectangular community garden. The length of the garden should be at least 110 ft, and the distance around should be no more than 380 ft. Which system of inequalities and graph represents the possible dimensions of the garden? I don't have the graph options for this, and I actually don't have any access to connexus stuff so idk the heck this is about L > or equal to 110 2(100) + 2W < or equal to 380 is all I can do for this one.

jigglypuff314 (jigglypuff314):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD per month to his collection while Nita is adding 5 CDs per month to her collection. After how many months will Tom and Nita have the same number of CDs? solve for x: 5x + 18 = x + 30

jigglypuff314 (jigglypuff314):

Kendra owns a restaurant. She charges $3.00 for two eggs and one piece of toast and $1.80 for one egg and one piece of toast. How much does Kendra charge per egg? How much per piece of toast? system of equations: 3e + t = 3 e + t = 1.8 elimination method(B)

jigglypuff314 (jigglypuff314):

y = 5x + 4 y = 5x – 3 I don't have graphs... they would be NON-intersecting parallel lines one of the lines has y-intercept 4 while the other line has y-intercept -3

jigglypuff314 (jigglypuff314):

y = –5x – 2 y + 2 = –5x I don't have graphs. they would be overlapping lines.

jigglypuff314 (jigglypuff314):

2x – y = –7 4x – y = –4 elimination method so you get 2x = 3 solve for x plug back into original to get y

jigglypuff314 (jigglypuff314):

I'm fried. good night.